OpenStudy (anonymous):

a right triangle has a hypotenuse of length 39 and a leg of length 15 what is the length of the other leg? if necessary round your answer to two decimal places.

OpenStudy (anonymous):

are you familiar with Pythagoras' Theorem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a^2 + b^2 = c^2 a = 15 c = 39 b is your answer

OpenStudy (anonymous):

\[b = \sqrt{39^{2} - 15^{2}}\]

OpenStudy (anonymous):

b= 36

OpenStudy (anonymous):

yep